A certain organic substance used as a solvent in many reactions contains carbon, hydrogen, oxygen and sulphur. Weight % of hydrogen in the compound is 7.7. The weight ratio C:O:S = 3:2:4, What is the least possible molar mass (in g) of the compound ?

Grade:12th pass

2 Answers

Vikas TU
7 years ago
Let x be the no. of H atoms in the compund then,
{ (1*x)/(1*x + 3*12 + 2*16 + 4*32) } * 100 = 7.7
solving for x we get,
x = 1509.2/92.3 = > 16.35
Total molecular mass = > 16.35*1 + 3*12 + 2*16 + 4*32 = > 212.35 g
Anurag Singh
5 years ago
Anurag singh
 
Hello friends, 
Mass of %hydrogen=7.7   simplest ratio =6
Mass of %carbon=30.4.    Simplest ratio=2
Mass of %oxygen=20.52.   Simplest ratio =1
Mass of %sulphur =41.04.    Simplest ratio =1
 
 The empirical formula is C2H6OS by going by the steps the simplest ratio will help you
 
The molar mass of the compound =2*12+6*1+16+32
=78
Here is the answer .good luck for exams.....